# Investigating pump harmonics generation in a SNAIL-based Traveling Wave Parametric Amplifier

### v2

@@ 2 Device description
The Reversed Kerr TWPA under investigation is nominally identical to the one reported in [7] and it is sketched in Fig. 1 (a). In the figure, $I_{C}$ indicates the critical current, $C_{g}$ indicates the unit cell capacitance to ground, $C_{J}$ the unit cell Josephson capacitance and $L$ the inductance per unit cell. Additional details on circuit parameters are reported in Appendix. A key design aspect for this device is that SNAILs are oriented in such a way that the external magnetic flux has opposite sign for adjacent unit cells, aiming at the suppression of the overall 3WM nonlinearity [32] and therefore of SHG.
@@
Figure 2: Couple Mode Equation (CME) theory predictions for gain and (a) THG of the pump as a function of pump power and external magnetic flux. Pump Pump frequency $f_{p}=4\,\text{GHz}$, signal frequency $f_{s}=3.9\,\text{GHz}$. The plots have been obtained with the open source software tool CMETANA (Couple Mode Equation Theory Analyzer) [35] which provides a user-friendly environment for CME predictions of JTWPAs starting from unit cells parameters. The exact analytical expressions are reported in Appendix and in CMETANA’s documentation.
@@ 3 Experimental results
In Fig. 3 (a), we report the device transmission ($S_{21}$ parameter) measured with a Vector Network Analyzer (VNA) with and without the presence of the pump tone at frequency $f_{p}=4\,\text{GHz}$, when external magnetic flux $\Phi_{\text{ext}}/\Phi_{0}$ is equal to 0.5. We deliberately chose the VNA settings in order to be able to detect the pump. When the pump is on, we observe 4WM gain along with the generation of the pump’s harmonics. To further investigate the generation of 4WM gain in the presence of pump harmonics, we perform a detailed spectral power characterization of the relevant frequency tones using a spectrum analyzer (SA). We set the pump frequency $f_{p}=4\,\text{GHz}$ and we measure the intensity of a signal tone, at frequency $f_{s}=3.9\,\text{GHz}$, as a function of the external magnetic flux and pump power (3 (b)). In the same conditions, we also measure the SHG, at frequency $f_{\text{SH}}=8\,\text{GHz}$, and the THG of the pump, at frequency $f_{\text{TH}}=12\,\text{GHz}$. The signal power at device input is $P_{s}=-110\,\text{dBm}$. The experimental results of harmonics measurements are reported in Fig. 4.
We observe 4WM gain (3 (b)) and THG (4 (b)) which periodically change with the applied magnetic flux, following the behaviour of the SNAIL 4WM nonlinear coefficient, as qualitatively expected from CME theory (Fig. 1 (b)). In contrast with what is ideally predicted for a SNAIL JTWPA with alternated flux polarity structure, we observe a non-negligible flux depended SHG power spectral density which for some flux values is comparable with the THG one. In addition, when the pump input power at the device exceeds roughly -71 dBm, we measure a sudden drop in the device’s experimental transmission due to the saturation of the Josephson junction critical current which is not described by CME theory.
@@ 4 Numerical simulations
Fabrication-related imperfections are also included in the numerical simulations. Specifically, we introduce a spread in the critical currents of all Josephson junctions by assuming a random Gaussian distribution with mean values equal to the ones estimated from fabrication parameters and linear experimental characterization of the device (see Appendix). The standard deviation of the random Gaussian distribution varies from $\delta/\braket{I}=0$ (ideal case in which all junctions with the same designed size have identical critical current) to $\delta/\braket{I}=15\%$ (typical upper limit for Josephson junctions’ arrays fabrication tolerance [30, 31]).
@@
Finally, an important outcome following from the numerical simulations is the influence of the spread of JJ’s critical currents on the generation of the second harmonic of the pump. While for an ideal device (no spread in the critical current values) SHG of the pump is found to be negligible, its intensity increases as the spread percentage increases, showing the best qualitative agreement with the flux-dependent experimental behaviour for typical values of fabrication-related JJ’s critical current spread of $5\%$. For a more accurate comparison with the experimental results, in Fig. 7, we report horizontal cuts of both experimentally and numerically obtained 2d-color-maps for a selected pump power of $-74\,\text{dBm}$, highlighting the simulations results for $0\%$ and $5\%$ spread.
@@ Acknowledgements
This work is supported by the European Union under Horizon Europe 2021-2027 Framework Programme Grant Agreement no. 101080152 and under Horizon 2020 Research and Innovation Programme Grant Agreement no. 731473 and 101017733; and by PNRR MUR project PE0000023-NQSTI.
@@ Appendix D Experiments in a large range of external magnetic flux
The observed slow modulation with the external flux can be explained by the fact that during the experiment no magnetic shielding has been adopted, making the measurements sensitive to additional environmental magnetic fields. Specifically, as recently reported in [38], the Josephson junction critical current can be affected by magnetic field lines along the insulating layer of Josephson junctions due to the ”bending” of the total perpendicular magnetic field.
@@ Appendix E Coupled mode equations theory
Appendix F SNAIL systematic fabrication imperfections influence

### v3

@@ 2 Device description
The Reversed Kerr TWPA under investigation is nominally identical to the one reported in [7] and it is sketched in Fig. 1 (a). In the figure, $I_{C}$ indicates the critical current, $C_{g}$ indicates the unit cell capacitance to ground, $C_{J}$ the unit cell Josephson capacitance and $L$ the inductance per unit cell. Additional details on circuit parameters are reported in Appendix. A key design aspect for this device is that SNAILs are oriented in such a way that the external magnetic flux has opposite sign for adjacent unit cells (alternating flux polarity), aiming at the suppression of the overall 3WM nonlinearity [32] and therefore of SHG.
@@
Figure 2: Couple Mode Equation (CME) theory predictions for gain (a) and THG of the pump (b) as a function of pump power and external magnetic flux. Pump Pump frequency $f_{p}=4\,\text{GHz}$, signal frequency $f_{s}=3.9\,\text{GHz}$. The signal power is assumed to be lower than 1 dB compression point power of the device. The plots have been obtained with the open source software tool CMETANA (Couple Mode Equation Theory Analyzer) [35] which provides a user-friendly environment for CME predictions of JTWPAs starting from unit cells parameters. The exact analytical expressions are reported in Appendix and in CMETANA’s documentation.
@@ 3 Experimental results
In Fig. 3 (a), we report the device transmission ($S_{21}$ parameter) measured with a Vector Network Analyzer (VNA) with and without the presence of the pump tone at frequency $f_{p}=4\,\text{GHz}$, when external magnetic flux $\Phi_{\text{ext}}/\Phi_{0}$ is equal to 0.5. We deliberately chose the VNA settings in order to be able to detect the pump. When the pump is on, we observe 4WM gain along with the generation of the pump’s harmonics. The VNA plots have an illustrative purpose, providing an example of the frequency-dependent behaviour of the device when pumped. To accurately investigate the generation of 4WM gain in the presence of pump harmonics, we perform a detailed spectral power characterization of the relevant frequency tones using a spectrum analyzer (SA). We set the pump frequency $f_{p}=4\,\text{GHz}$ and we measure the intensity of a signal tone, at frequency $f_{s}=3.9\,\text{GHz}$, as a function of the external magnetic flux and pump power (3 (b)). In the same conditions, we also measure the SHG, at frequency $f_{\text{SH}}=8\,\text{GHz}$, and the THG of the pump, at frequency $f_{\text{TH}}=12\,\text{GHz}$. The measured power is calibrated at the output of the device. The signal power at device input is $P_{s}=-110\,\text{dBm}$. The experimental results of harmonics measurements are reported in Fig. 4.
We observe 4WM gain (3 (b)) and THG (4 (b)) which periodically change with the applied magnetic flux, following the behaviour of the SNAIL 4WM nonlinear coefficient, as qualitatively expected from CME theory (Fig. 1 (b)). In contrast with what is ideally predicted for a SNAIL JTWPA with alternated flux polarity structure, we observe a non-negligible flux depended SHG power spectral density which for some flux and pump power values is comparable or larger than the THG one. We stress that for an accurate quantitative comparison between the measured spectral power of SHG and THG one should also consider that losses in the device under study increase for larger frequencies, specifically they approximately increase of 1 dB per GHz [7]. In addition, when the pump input power at the device exceeds roughly -71 dBm, we measure a sudden drop in the device’s experimental transmission due to the saturation of the Josephson junction critical current which is not described by CME theory.
@@ 4 Numerical simulations
Fabrication-related imperfections among different unit cells are also included in the numerical simulations. Specifically, we introduce a spread in the critical currents of all Josephson junctions in the full device (700 cells) by assuming a random Gaussian distribution with mean values equal to the ones estimated from fabrication parameters and linear experimental characterization of the device (see Appendix). The standard deviation of the random Gaussian distribution varies from $\delta/\braket{I}=0$ (ideal case in which all junctions with the same designed size have identical critical current) to $\delta/\braket{I}=15\%$ (typical upper limit for Josephson junctions’ arrays fabrication tolerance [30, 31]).
@@
Finally, an important outcome following from the numerical simulations is the influence of the spread of JJ’s critical currents on the generation of the second harmonic of the pump. While for an ideal device (no spread in the critical current values) SHG of the pump is almost fully suppressed for pump powers below the saturation of the critical current (about -70 dB), its intensity increases as the spread percentage increases, showing the best qualitative agreement with the flux-dependent experimental behaviour for typical values of fabrication-related JJ’s critical current spread of $5\%$.
We interpret the increase of the SHG with increasing spread in the Josephson junctions’ critical currents as an effect of the increasing disorder in the alternating flux polarity design. The latter ideally guarantees a full cancellation of the 3WM nonlinearity in the approximation of wavelengths much larger than the unit cells [32]. This is valid because the 3WM coefficient $\beta$ associated with each SNAIL (see Eq. 4 in Appendix) is an odd function of the external flux. In the case of 0 % spread, adjacent SNAILs are identical but with opposite flux polarity. This means that the corresponding $\beta$ coefficient will have identical absolute value but opposite sign, resulting in an overall suppression of 3WM processes like SHG. By switching on the spread in junction critical currents, adjacent SNAILs and the corresponding $\beta$ amplitude are not identical anymore, and their difference gets bigger as the spread in junction critical current increases, thus increasing the overall $\beta$ and so the efficiency of SHG. The residual SHG at 0% spread below the saturation power is attributed to a deviation from the assumption of fully cancelled 3WM nonlinearity for the alternated flux polarity design, corresponding to the approximation of wavelengths much larger than the unit cells [32].
For a more accurate comparison with the experimental results, in Fig. 7, we report horizontal cuts of both experimentally and numerically obtained 2d-color-maps for a selected pump power of $-74\,\text{dBm}$, highlighting the simulations results for $0\%$ and $5\%$ spread.
@@
In Fig 7, we observe qualitative agreement between experiments and theory for gain and SHG, where the experimental curve lies in between the simulated ones with 0% and 5% spread. However, for the THG case, the simulations are both overestimating the power observed experimentally. We attribute this discrepancy to the fact that our model does not include losses, which in the real device are considerably higher at the frequency of the THG than at the SHG and signal frequencies [7]. The inclusion of losses in the simulations is a promising perspective for future investigations in order to improve the agreement between simulations and experiments at higher frequencies.
@@ Acknowledgements
This work is supported by the European Union under Horizon Europe 2021-2027 Framework Programme Grant Agreement no. 101080152 and under Horizon 2020 Research and Innovation Programme Grant Agreement no. 731473 and 101017733; and by PNRR MUR project PE0000023-NQSTI. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 899561.
@@ Appendix D Experiments in a large range of external magnetic flux
The observed slow modulation with the external flux can be explained by the fact that during the experiment no magnetic shielding has been adopted, making the measurements sensitive to additional environmental magnetic fields. Specifically, as recently reported in [38, 39], the Josephson junction critical current can be affected by magnetic field lines along the insulating layer of Josephson junctions due to the ”bending” of the total perpendicular magnetic field.
@@ Appendix E Coupled mode equations theory
Appendix F Frequency-dependent gain profile
In Fig. 12 we report a comparison between the experimental frequency-dependent gain profile (differential trace of the “pump ON” and “pump OFF” traces in Fig. 3 (a)) and simulations for different spread values of the Josephson junctions critical current.
Figure 12: Comparison between the experimental frequency dependent gain profile (differential trace of the “pump ON” and “pump OFF” traces reported in Fig. 3 (a)) and WRSpice simulations for spread values $\delta=0\%$, $\delta=5\%$ and $\delta=10\%$. External magnetic flux $\Phi_{\text{ext}}/\Phi_{0}=0.5$. Pump power at the device is $P_{p}=-74\,\text{dBm}$.
[…]
Appendix G SNAIL systematic fabrication imperfections influence